OpenStudy (anonymous):

the volume of a spherical balloon is increasing at a rate of 4 cubic meters per minute. How fast is the surface area increasing when the radius is 3 meters

OpenStudy (bahrom7893):

dv/dt = 4 r = 3, da/dt - ?

OpenStudy (bahrom7893):

a = 4pi*r^2 da/dt = 8pi*r*(dr/dt) = 24pi(dr/dt)

OpenStudy (bahrom7893):

dv/dt = 4pi*r^2(dr/dt) 1 = 9pi(dr/dt) dr/dt = 1/(9pi)

OpenStudy (bahrom7893):

da/dt = 24pi/(9pi) = 24/9 = 8/3

OpenStudy (bahrom7893):

check for arithmetic errors

OpenStudy (anonymous):

hey is there anyway you can rewrite your second response out again i cant figure out how to write it out. thanks